Mitch is a kicker in the Superbowl. He must make a field goal in order to win the game. He is 35 yards away from the goal posts when he kicks the football with a velocity of 19 m/s at an angle of 48
Answer #1


Range R = u^2*sin(2*48)/g = 19^2*sin(48)/9.8   = 27.37 m

but 1 m = 1.093 yards = 3.28 foot
then arnge R = 27.37*1.093 = 29.9 Yards and 27.37*3.28 = 89.77 feets which is greater than 10 ft......
A) answer for A) is Yes

B)T = u*sin(48)/g = 19*sin(48)/9.8 = 1.44 sec...

C) in 1.44 sec ..it travels horizontally a distance of (19*cos(48)*1.44) = 18.3 m = 18.3/3.28 = 5.58 ft....

after 5.58 ft the ball starts falling and goal post is at 10 ft distance so when it gets gaol post it was falling